OpenStudy (anonymous):

Find Roots : 4x^3+16x^2-9x-36=0 sum of two of the roots =0.

OpenStudy (anonymous):

Use the factor theorem to find the first root

OpenStudy (anonymous):

pls use sum and product of roots properties.

OpenStudy (anonymous):

let the two roots be \[\alpha &\]

OpenStudy (anonymous):

Roots : p,q,r p+q+r = -b/a pq+qr+rp= c/a pqr = -d/a

OpenStudy (anonymous):

ok so a+b+c = -16/4=-4

OpenStudy (anonymous):

as a+b=o c= -4

OpenStudy (anonymous):

nw (+4) is a factor of the given polynomial ryt

OpenStudy (anonymous):

i mean (x+4)

OpenStudy (anonymous):

no factors

OpenStudy (anonymous):

i want it only from the sum and product

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

thanks for your efforts.

OpenStudy (anonymous):

nw if u divide the given polynomial by (x+4) u will get a quadraitic equation de if u split de middle term u will get the odr roots

OpenStudy (anonymous):

ok i tried to my stndrd :)

OpenStudy (anonymous):

In which class are you studying?

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

just passed 10th

OpenStudy (anonymous):

so intrmediate 1st yr?

OpenStudy (anonymous):

cant say now everything is new now

OpenStudy (anonymous):

y so?

OpenStudy (anonymous):

course seems to be very different.

OpenStudy (anonymous):

that you have studied in 9th, was just an intorduction. In 11 its full detailed.

OpenStudy (anonymous):

what result you got?

OpenStudy (anonymous):

we'll get it only aftr 10th, cce ryt!